# Sparse R-CNN OBB: Ship Target Detection in SAR Images Based on Oriented Sparse Proposals

### v1

@@ Abstract
We present Sparse R-CNN OBB, a novel framework for the detection of oriented objects in SAR images leveraging sparse learnable proposals. The Sparse R-CNN OBB has streamlined architecture and ease of training as it utilizes a sparse set of 300 proposals instead of training a proposals generator on hundreds of thousands of anchors. To the best of our knowledge, Sparse R-CNN OBB is the first to adopt the concept of sparse learnable proposals for the detection of oriented objects, as well as for the detection of ships in Synthetic Aperture Radar (SAR) images. The detection head of the baseline model, Sparse R-CNN, is re-designed to enable the model to capture object orientation. We also fine-tune the model on RSDD-SAR dataset and provide a performance comparison to state-of-the-art models. Experimental results shows that Sparse R-CNN OBB achieves outstanding performance, surpassing other models on both inshore and offshore scenarios. The code is available at: www.github.com/ka-mirul/Sparse-R-CNN-OBB*.
Index Terms:
Convolutional neural network, deep learning, oriented ship detection, sparse learnable proposals, synthetic aperture radar (SAR).
I Introduction
@@
The ship detection task involves localizing ships within an image and determining their classes. In traditional SAR ship detection systems, the process generally involves three key stages: pre-processing, candidate extraction, and discrimination. The pre-processing step may involve enhancing the quality of the input image, reducing noise, and conducting land-sea segmentation to avoid extracting non-ocean targets. In candidate extraction stage, potential ship targets are identified. Methods based on the Constant False Alarm Rate (CFAR) technique are commonly used for this task. CFAR-based detectors are adaptive threshold systems that estimate sea clutter statistics surrounding a potential target using an assumed background probability density function. This estimation enables them to maintain a consistent, acceptable probability of false alarm (PFA) [1, 2]. Finally, the discrimination stage ensures that only genuine ships are retained for further analysis. Although traditional detectors generally deliver satisfactory performance, their accuracy is prone to variability due to their heavy reliance on hand-crafted parameters assigned at each stage. Furthermore, the lack of end-to-end design in these approaches increases design complexity, making both training and adaptation to other datasets more challenging.
@@
The remainder of this article is organized as follows, Section II explains the concept of sparse learnable proposals. Section III provides the implementation details of Sparse R-CNN OBB . Experimental details are provided in Section IV, while the experimental results and performance comparison to other models are covered in Section V. Finally, Section VI provides concluding remarks.
II Related Work
II-A Sparse Learnable Proposals
The key idea of sparse learnable proposals, first introduced in Sparse R-CNN [12], is the substitution of hundreds of thousands of box candidates (anchors) with a small set of proposal boxes. From a technical standpoint, learnable proposals represent an improved version of traditional proposals, enhanced with additional learnable features (256-dimensional in this instance), referred to as proposal features. This addition allows each proposal to embed not only box-related parameters but also unique 256-dimensional instance characteristics.
@@ III Sparse R-CNN OBB
III-A Backbone
Sparse R-CNN OBB employs a ResNet-50 backbone with a Feature Pyramid Network (FPN) for multi-scale feature fusion, referred to as ResNet-50-FPN. Initially, residual blocks of ResNet-50 (res2, res3, res4, res5) output feature maps C2, C3, C4, and C5, which are then passed to the FPN. Afterwards, to standardize their channel dimensions (256, 512, 1024, 2048), a 1×1 lateral convolution is applied to align all these maps to 256 channels. The aligned maps are further iteratively fused in a top-down manner (see Fig.1), with smaller maps upsampled to match the larger ones. Finally, a final 3×3 convolution refines the fused maps, producing outputs P2, P3, P4, and P5.
III-B Rotated Sparse Learnable Proposals
We exploit the concept of sparse learnable proposals, originally built for detecting non-rotated objects, to further accommodate the detection of oriented objects. In our pipeline, each proposal contains 256-dimensional learnable proposal features and 5-dimensional box parameters $x_{p}$, $y_{p}$, $w_{p}$, $h_{p}$, $\theta_{p}$. The ($x_{p}$, $y_{p}$) represents the center point of the proposal while ($w_{p}$, $h_{p}$, $\theta_{p}$) respectively represent the width, height, and orientation of the proposal.
Figure 1: Flowchart of the proposed Sparse R-CNN OBB.
@@ III-C Loss Function
In this formulation, the total loss is a weighted sum of the focal loss $\mathcal{L}_{\text{cls}}$, L1 loss $\mathcal{L}_{\text{iou}}$, and IoU loss $\mathcal{L}_{\text{iou}}$, with coefficients $\lambda_{\text{cls}},\lambda_{\text{L1}},\lambda_{\text{iou}}$. The training loss, computed only for matched pairs, follows the same structure as the matching loss and is normalized by the number of objects in the batch.
IV Experiment Setup
IV-A Dataset
@@ IV-B Evaluation
IV-C Hyperparameters and Environment
Sparse R-CNN OBB is set up with weights pre-trained on ImageNet [16]. The model is then trained using the Adam optimizer [17] with a base learning rate of $7.5\times 10^{-5}$ for 150 epochs, with momentum set to 0.9, weight decay set to $1\times 10^{-4}$. A warm-up strategy is implemented for the first 1000 iterations, with a learning rate 1% of the base rate. The base learning rate is further reduced by a factor of 10 at the 130th and 140th epochs. Training uses a batch size of 8 on two NVIDIA RTX 2080 GPUs with Detectron2 [18] in PyTorch on Ubuntu 22.04.
V Results and Discussion
@@
V-A Effect of Number of Proposals
@@
TABLE I: Experiment on Number of Proposals
Proposals AP50(%) Model size (M) Training time (h) FPS
@@
According to Table I, increasing the number of proposals marginally enhances performance, but at the cost of longer training times and reduced inference speed. The results also demonstrate that using more proposals does not significantly increase model size. This is because the increase in proposals only adds parameters for storing the 5-dimensional proposal boxes and 256-dimensional proposal features, which are minimal compared to the overall model size.
@@
Figure 2: Detection results across different environments, with predictions (magenta) and ground truth (green). The first two left columns show results in inshore regions, while the remaining columns correspond to offshore regions.
V-B Comparison to SOTA Models
In this section, we compare the performance of Sparse R-CNN OBB against state-of-the-art algorithms. As shown in Table II, Sparse R-CNN OBB outperforms other methods across all metrics, except for $AP_{50}Inshore$. In mixed-scene test sets, reflected by $AP_{50}$, our method demonstrates superior performance, surpassing all state-of-the-art algorithms by a significant margin ranging from 2.47% to 25.12%, with CFA [9] being the closest competitor. Compared to other two-stage detectors, our model demonstrates a superior performance with an improvement of 2.94% over Oriented R-CNN [7], the top-performing algorithm in this category. This performance comparison shows that Sparse R-CNN OBB is robust for detecting ships in typical operational ocean environments, where both inshore and offshore backgrounds may coexist.
TABLE II: Performance Comparison on RSDD-SAR Dataset.
@@ V-B Comparison to SOTA Models
In inshore scenes, the Sparse R-CNN OBB surpasses other SOTA algorithms, with the exception of CFA. However, the performance margin is minimal ($<$0.02%), positioning the method as the second highest. Despite this, our proposed model remains highly competitive and can be considered as a top-performing method. This result indicates that Sparse R-CNN OBB can better detect ships in complex environments, whereas other methods struggle with higher false detection rates.
Lastly, for offshore scenes, Sparse R-CNN OBB demonstrates exceptional performance, achieving more than 96% accuracy and surpassing other models by at least a 1.87% margin. This results signifies that Sparse R-CNN OBB is more robust against noisy SAR ocean images, which may also be interfered by sea waves and ship-generated wakes.
@@
VI Conclusion
In this article, we proposed Sparse R-CNN OBB, a new pipeline for oriented ship detection in SAR imageries based on sparse learnable proposals. The adoption of sparse proposals concept ensures a simple design for our model. Furthermore, the incorporation of rich instance features from learnable proposals improves the model performance.  Sparse R-CNN OBB achieves superior performance, surpassing all state-of-the-art benchmarks on RSDD-SAR dataset by a margin of at least 2.13%.

### v2

# Sparse R-CNN OBB: Ship Target Detection in SAR Images Based on Oriented Sparse Learnable Proposals

@@ Abstract
We present Sparse R-CNN OBB, a novel framework for the detection of oriented objects in SAR images leveraging sparse learnable proposals. The Sparse R-CNN OBB has streamlined architecture and ease of training as it utilizes a sparse set of 300 proposals instead of training a proposals generator on hundreds of thousands of anchors. To the best of our knowledge, Sparse R-CNN OBB is the first to adopt the concept of sparse learnable proposals for the detection of oriented objects, as well as for the detection of ships in Synthetic Aperture Radar (SAR) images. The detection head of the baseline model, Sparse R-CNN, is re-designed to enable the model to capture object orientation. We train the model on RSDD-SAR dataset and provide a performance comparison to state-of-the-art models. Experimental results show that Sparse R-CNN OBB achieves outstanding performance, surpassing most models on both inshore and offshore scenarios. The code is available at: www.github.com/ka-mirul/Sparse-R-CNN-OBB.
Index Terms—  Convolutional neural network, deep learning, oriented ship detection, sparse learnable proposals, synthetic aperture radar (SAR).
1 Introduction
@@
The ship detection task involves localizing ships within an image and determining their classes. In traditional SAR ship detection systems, the process generally involves three key stages: pre-processing, candidate extraction, and discrimination. The pre-processing step may involve enhancing the quality of the input image, reducing noise, and conducting land-sea segmentation to avoid extracting non-ocean targets. In candidate extraction stage, potential ship targets are identified. Methods based on the Constant False Alarm Rate (CFAR) technique are commonly used for this task. CFAR-based detectors are adaptive threshold systems that estimate sea clutter statistics surrounding a potential target using an assumed background probability density function. This estimation enables them to maintain a consistent, acceptable probability of false alarm (PFA) [1]. Finally, the discrimination stage ensures that only genuine ships are retained for further analysis. Although traditional detectors generally deliver satisfactory performance, their accuracy is prone to variability due to their heavy reliance on hand-crafted parameters assigned at each stage. Furthermore, the lack of end-to-end design in these approaches increases design complexity, making both training and adaptation to other datasets more challenging.
@@
The remainder of this article is organized as follows, Section 2 explains the concept of sparse learnable proposals. Section 3 provides the implementation details of Sparse R-CNN OBB . Experimental details are provided in Section 4, while the experimental results and performance comparison to other models are covered in Section 5. Finally, Section 6 provides concluding remarks.
2 Related Work
2.1 Sparse Learnable Proposals
The concept of utilizing sparse proposal detectors has been explored in prior research [11, 12, 13, 14], with G-CNN [11] serving as a pioneering approach in this domain. Instead of relying on a proposal generator, G-CNN initiates with a multi-scale grid comprising 180 fixed bounding boxes. These initial boxes are iteratively refined by a regressor that adjusts their position and scale to better align with the objects in the image. At the time of its introduction, G-CNN achieved performance comparable to Fast R-CNN [15], which operates with approximately 2,000 bounding boxes. However, G-CNN lags behind the performance of its successor, Faster R-CNN [16].
DETR [12] advanced the field by proposing a transformer-based architecture that redefines object detection as a set prediction problem. DETR employs a fixed set of 100 learned object queries to simultaneously capture object relationships and global image context, enabling direct parallel predictions. Despite its conceptual elegance, DETR relies on dense interactions between object queries and global image features, resulting in slow convergence during training and hindering the development of a fully sparse detection pipeline
Sparse R-CNN [13] is the first framework to introduce a fully sparse proposal paradigm for object detection. Instead of relying on extensive candidate boxes or dense interactions with global image features, Sparse R-CNN employs a compact set of learnable proposals, known as sparse learnable proposals, comprising both bounding box coordinates and feature embeddings. These proposals interact directly with Region of Interest (RoI) features in a one-to-one manner, eliminating the need for global context processing. This fully sparse interaction significantly reduces computational complexity while achieving efficient and precise object detection. The sparse learnable proposals in Sparse R-CNN provide key advantages over traditional dense proposals. They eliminate the need for a region proposal network (RPN) during training and non-maximum suppression (NMS) during inference as densely-overlapping predictions are inherently avoided. This results in a simplified architecture comprising only a backbone feature extractor, an interaction module, and classification/regression heads
The Sparse Anchoring Network (SAN) proposed in [14] represents another effort to incorporate sparse proposals by generating a reduced set of region proposals. However, SAN operates as a region proposal network (RPN) within the R-CNN framework, preventing it from achieving the simplicity and efficiency of the fully sparse design of Sparse R-CNN.
@@ 3 Sparse R-CNN OBB
3.1 Backbone
Sparse R-CNN OBB employs a ResNet-50 backbone with a Feature Pyramid Network (FPN) for multi-scale feature fusion, referred to as ResNet-50-FPN. As depicted in Fig. 2, residual blocks of ResNet-50 (res2, res3, res4, res5) first output feature maps C2, C3, C4, and C5, which are then passed to the FPN. Afterwards, to standardize their channel dimensions (256, 512, 1024, 2048), a 1×1 lateral convolution is applied to align all these maps to 256 channels. The aligned maps are further iteratively fused in a top-down manner, with smaller maps upsampled to match the larger ones. Finally, a final 3×3 convolution refines the fused maps, producing outputs P2, P3, P4, and P5.
3.2 Rotated Sparse Learnable Proposals
We exploit the concept of sparse learnable proposals, originally built for detecting non-rotated objects, to further accommodate the detection of oriented objects. In our pipeline, each proposal contains 256-dimensional learnable proposal features and 5-dimensional box parameters $x_{p}$, $y_{p}$, $w_{p}$, $h_{p}$, $\theta_{p}$. As depicted in Fig. 1, the ($x_{p}$, $y_{p}$) represents the center point of the proposal while ($w_{p}$, $h_{p}$, $\theta_{p}$) respectively represent the width, height, and orientation of the proposal.
Fig. 1: Oriented bounding box representation used in the proposed Sparse R-CNN OBB.
Fig. 2: Flowchart of the proposed Sparse R-CNN OBB.
@@ 3.3 Loss Function
In this formulation, the total loss is a weighted sum of the focal loss $\mathcal{L}_{\text{cls}}$, L1 bounding box regression loss $\mathcal{L}_{\text{iou}}$, and IoU loss $\mathcal{L}_{\text{iou}}$, with coefficients $\lambda_{\text{cls}},\lambda_{\text{L1}},\lambda_{\text{iou}}$. The $\lambda_{\text{cls}}$, $\lambda_{\text{L1}}$, and $\lambda_{\text{IoU}}$ coefficients were set to 2.0, 5.0, and 2.0, respectively, following the baseline Sparse R-CNN settings [13]. The training loss, computed only for matched pairs, follows the same structure as the matching loss and is normalized by the number of objects in the batch.
4 Experiment Setup
4.1 Dataset
@@ 4.2 Evaluation
4.3 Hyperparameters and Environment
Sparse R-CNN OBB is set up with weights pre-trained on ImageNet [19]. The model is then trained using the AdamW optimizer [20] with a base learning rate of $7.5\times 10^{-5}$ for 150 epochs, with momentum set to 0.9, weight decay set to $1\times 10^{-4}$. A warm-up strategy is implemented for the first 1000 iterations, with a learning rate 1% of the base rate. The base learning rate is further reduced by a factor of 10 at the 130th and 140th epochs. Training uses a batch size of 8 on two NVIDIA RTX 2080 GPUs with Detectron2 [21] in PyTorch on Ubuntu 22.04, hosted by the Advanced Computing Research Centre’s HPC systems at the University of Bristol.
5 Results and Discussion
@@
5.1 Effect of Number of Proposals
@@
Table 1: Experiment on Number of Proposals
Proposals AP50(%) Model size (M) Training Duration (h) FPS
@@
According to Table 1, increasing the number of proposals marginally enhances performance, but at the cost of longer training times and reduced inference speed. The results also demonstrate that using more proposals does not significantly increase model size. This is because the increase in proposals only adds parameters for storing the 5-dimensional proposal boxes and 256-dimensional proposal features, which are minimal compared to the overall model size.
@@
5.2 Comparison to SOTA Models
In this section, we compare the performance of Sparse R-CNN OBB against state-of-the-art algorithms. As shown in Table 2, Sparse R-CNN OBB outperforms other methods across all metrics, except for $AP_{50}Inshore$. In mixed-scene test sets, reflected by $AP_{50}$, our method demonstrates superior performance, surpassing all state-of-the-art algorithms by a significant margin ranging from 2.51% to 25.16%, with CFA [8] being the closest competitor. Compared to other two-stage detectors, our model demonstrates a superior performance with an improvement of 2.98% over Oriented R-CNN [6], the top-performing algorithm in this category. This performance comparison shows that Sparse R-CNN OBB is robust for detecting ships in typical operational ocean environments, where both inshore and offshore backgrounds may coexist.
Fig. 3: Detection results across different environments, with predictions (magenta) and ground truth (green). The first two left columns show results in inshore regions, while the remaining columns correspond to offshore regions.
Table 2: Performance Comparison on RSDD-SAR Dataset
@@ 5.2 Comparison to SOTA Models
In inshore scenes, Sparse R-CNN OBB achieves performance comparable to CFA, with a marginal gap of less than 0.02%. While CFA is tailored for densely packed, arbitrarily oriented objects using convex-hull-based adaptation, Sparse R-CNN OBB is specifically designed for SAR ship detection across both inshore and offshore settings. These results suggest that incorporating inshore-specific refinements could further enhance performance. Nonetheless, Sparse R-CNN OBB remains highly competitive, exhibiting strong robustness and lower false detection rates in complex environments. This result indicates that Sparse R-CNN OBB can better detect ships in complex environments, whereas other methods struggle with higher false detection rates.
Lastly, for offshore scenes, Sparse R-CNN OBB demonstrates exceptional performance, achieving more than 96% accuracy and surpassing other models by at least a 1.91% margin. This results signifies that Sparse R-CNN OBB is more robust against noisy SAR ocean images, which may also be interfered by sea waves and ship-generated wakes.
@@
6 Conclusion
In this article, we proposed Sparse R-CNN OBB, a new pipeline for oriented ship detection in SAR imageries based on sparse learnable proposals. The adoption of sparse proposals concept ensures a simple design for our model. Furthermore, the incorporation of rich instance features from learnable proposals improves the model performance.  Sparse R-CNN OBB achieves superior performance, surpassing all state-of-the-art benchmarks on the RSDD-SAR dataset by at least 2.51% in $AP_{50}$.